Circular No. 7448

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000ct IN ANONYMOUS GALAXY T. Puckett, Mountain Town, GA; A. Sehgal, Woodinville, WA; and

J. Newton, Osoyoos, BC, report their discovery of an apparent supernova (mag 17.5) on an unfiltered CCD frame (limiting mag 20.5) taken with the Puckett Observatory 0.60-m automated supernova patrol telescope on July 4.13 UT. The candidate is located at α = 17^{h}01^{m}03^{s}.64, δ = +33^{o}28′45′′.0 (equinox 2000.0), which is 7′′.0 west and 13′′.4 north of the center of the host galaxy. The new object was also present on an unfiltered CCD frame taken with the 0.40-m reflector at the Newton Observatory B & B on July 6.55 UT. It does not appear on Palomar Sky Survey images taken on 1993 Apr. 4, 1989 June 29 and 1954 May 27 (limiting mag ∼ 21.0, 21.0 and 20.0, respectively.

SUPERNOVA 2000cs IN MCG +07-34-15 S. Jha, P. Challis and R. Kirshner, Harvard-Smithsonian Center for

Astrophysics, report that spectra of SN 2000cs, taken with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph) by M. Calkins on July 2.23, 3.26 and 4.30 UT, and by A. Mahdavi on July 5.28, show it to be an unusual type II supernova near maximum light. The spectra exhibit a slightly red continuum with a broad P-Cyg Hα profile. The Hα emission FWHM is 11 800 km/s, and the expansion velocity measured to the Hα trough is 20 000 km/s, adopting the NASA/IPAC Extragalactic Database recession velocity of 10 596 km/s for the host galaxy. Other well-developed Balmer and He lines are present, with expansion velocities 11 400 km/s (He I, rest 587.6 nm), 9600 km/s (Hβ) and 8900 km/s (Hγ). See http://cfa-

www.harvard.edu/cfa/oir/Research/supernova/RecentSN.html.

A. V. Filippenko, R. Chornock and M. Modjaz, University of California at Berkeley, report that an uncalibrated spectrum with the Lick Observatory’s Shane 3-m reflector on July 6 suggests that this is a somewhat peculiar type II supernova, dominated by a strong, broad, Hα emission line, with weaker Na D (or He I) absorption and some other features.

V1494 AQUILAE R. Novák, Copernicus Observatory; A. Retter, Keele University; L.

Cook, Concord, CA; and D. Korčaková and J. Janík, Masaryk University, write: “CCD photometry of V1494 Aql (Nova Aql 1999 No. 2) through an R filter on June 7, 8, 9, 10, 15 and 16 reveals a sinusoidal variation with period 0.0627 ± 0.0001 day and full amplitude 0.03 mag.”

2000 July 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Brian G. Marsden